# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v3

@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $B_{h,v}(1)_{2}$ exists and $c=\sum_{n\geq 1}|b_{n}^{h}|<1$. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. Let $B_{h,v}(x)\neq B_{h}(x)$ and $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)\neq B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
@@
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. Proposition 2.15 gives the stated formula:
$\lim_{n\to\infty}a_{n}^{h,v}={\textstyle C_{h,\,v}(1)_{1}=C_{h,\,v}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\big)\,.}$
@@
Theorem 4.9 (12: $v\neq 1$, con, $U(1)_{1}=+\infty$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive convex light weight such that $v\in V(h)$. If $B_{h}(1)_{1}=+\infty$ then
$\lim_{n\to\infty}a_{n}^{h,v}=1\,.$
Proof. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-x}$. Hence $0\leq b_{n}^{h},b_{n}^{h,v},c_{n}^{h,v},e_{n}^{h,v}\leq d_{n}^{h}=1$ for every $n\in\mathbb{N}_{0}$ and we can use continuous and non-decreasing functions
$F_{B_{h}},\,F_{B_{h,v}},\,F_{C_{h,v}},\,F_{E_{h,v}}\colon[0,\,1)\to[0,\,+\infty)\,.$
Any shortest walk in $G(h)$ joining $1$ and $v$ shows that for some $m\in\mathbb{N}$ we have $e_{m}^{h,v}>0$. Thus the function $F_{E_{h,v}}(x)$ is on the interval $(0,1)$ positive and increasing. We deduce from it that $F_{B_{h,v}}(x)$ is bounded on $[0,1)$.
Suppose for the contrary that $\lim_{x\to 1}F_{B_{h,v}}(x)=+\infty$. By item 3 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality $F_{C_{h,v}}(x)=F_{B_{h,v}}(x)F_{E_{h,v}}(x)$. Thus $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$. Since $F_{C_{h,v}}(0)=0$, $F_{C_{h,v}}(x)$ increases and is continuous and $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$, we obtain a (unique) number $x_{0}\in(0,1)$ such that $F_{C_{h,v}}(x)<1$ for $x\in[0,x_{0})$ and $\lim_{x\to x_{0}^{-}}F_{C_{h,v}}(x)=1$. By item 2 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality
$F_{B_{h}}(x)=F_{B_{h,v}}(x)+F_{C_{h,v}}(x)^{2}F_{B_{h}}(x)\,\,.$
Hence for every $x\in[0,x_{0})$ we have $F_{B_{h}}(x)=\frac{F_{B_{h,v}}(x)}{1-F_{C_{h,v}}(x)^{2}}$. Since $F_{B_{h,v}}(0)=1$, it follows that $\lim_{x\to x_{0}^{-}}F_{B_{h}}(x)=+\infty$. But this contradicts the continuity of $F_{B_{h}}(x)$ at $x_{0}$.
Using item 2 of Proposition 4.1, the boundedness of $F_{B_{h,v}}(x)$ and Proposition 2.23, we get $C_{h,v}(1)_{1}=1$. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}=1$. $\Box$
@@ 5 Concluding remarks
In [17] we consider the same problem of determination of formulas for the quantities
$A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))$}
for formal weights $h$. These are maps
@@
The main novelty of our generalization is this. In an equality like
@@
in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain. Construction of this probability space is far from trivial; see, for example, [6]. In our approach, the right-hand side is completely self-contained.

### v6

@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
1. Suppose that $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
and that $(B_{h,v})_{1/\xi}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}}\big)\,.$
2. Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
and that $B_{h,v}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
then $(B_{h})(1)_{2}\neq 0$ and
$A_{h,v}(1)_{2}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
@@
1. Let $h$ be $\xi$-converx with $\xi\in\mathbb{C}_{1}$, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives the stated formula:
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}={\textstyle(C_{h,\,v})_{1/\xi}(1)_{1}=(C_{h,\,v})_{1/\xi}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\big)\,.}$
2. Let $h$ be 0-convex, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. The beginning of the previous case gives for $\xi=1$ that $\big(C_{h,v}(1)_{2}\big)^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)$, the stated formula follows. $\Box$
Theorem 4.10 (12: $v\neq 1$, con, $U(1)_{1}=\infty$)
1. If $\xi\in\mathbb{C}_{1}$, $v$ is in $\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=1\,.$
2. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, $B_{h},C_{h,v}\in H_{1}$, $(B_{h,v})_{1/\xi}(1)_{1}$ exists and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
Proof. 1. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. By the subseries bound principle we have
$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}b_{n}^{h,v},\,\xi^{-n}c_{n}^{h,v},\,\xi^{-n}e_{n}^{h,v}\leq\xi^{-n}d_{n}^{h}=1$
for every $n\in\mathbb{N}_{0}$. Hence $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$ and $E_{h,v}(x)$ are in $H_{1}$. Any shortest walk joining $1$ and $v$ in $G(h)$ shows that for some $m\in\mathbb{N}$ we have $\xi^{-m}e_{m}^{h,v}>0$. Thus the function $F_{(E_{h,v})_{1/\xi}}(x)$ is positive and increasing on $(0,1)$. Since $F_{(C_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ by item 2 of Proposition 4.1, $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded by item 3. Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$, Proposition 2.15 gives
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=(C_{h,v})_{1/\xi}(1)_{2}\,.$
By item 2 of Proposition 4.1, Propositions 2.2, 2.4 and 2.5, and by Theorem 2.18 we have
$(C_{h,v}^{2})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1\,$,}
because $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ and $\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)=B_{h}(1)_{1}=+\infty$. By Proposition 2.25, $(C_{h,v})_{1/\xi}(1)_{2}$ exists and
$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=\big((C_{h,v})_{1/\xi}\big)^{2}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{1}=1\,.$
Since $(C_{h,v})_{1/\xi}(1)_{2}\geq 0$, we have $(C_{h,v})_{1/\xi}(1)_{2}=1$ and the stated formula follows.
2. By item 1 of Proposition 4.1 and by Proposition 2.31, $A_{h,v}(x)=C_{h,v}(x)$. By item 2 of Proposition 4.1 and by Theorems 2.16 and 2.17,
$\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{\pm\infty}=1\,.$}
By the continuity of the functions $F_{\dots}$,
$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)=\lim_{x\nearrow 1}F_{(C_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
@@ 5 Concluding remarks
In [17] we consider the same problem of determining formulas for quantities such as $A_{h,v}(1)=\sum_{n\geq 0}h(W(v,n))$ in the case of formal weights $h$. These are maps
@@
The main novelty of our generalization is this. In equalities like
@@
that was stated in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain. Construction of this probability space is far from trivial; see, for example, [6]. In our generalization, the right-hand side is completely self-contained.